the value of lim x tends to a cos x – cos a / cot x – cot a is ….................
 

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Here by putting the value a in place of x we get that  L Hospital rule can be applied here.
              By applying this we get,
             Lim (xàa) Cos x- Cot x / Cot x – Cot a
            = lim (xàa) -Sinx/ - Cosec^2 x
           = lim (xàa) Sin^3 x
            = Sin^3 a.
Hence the limit values is sin^3a.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)